1. On 12/31, Choco acquired all assets and liabilities of Cake by issuing 40,000 shares of its common stock when the market value (=fair value) is $32/share and this combination is a statutory merger (Cake was dissolved). Choco has common stock with $15 par, 50,000 shares outstanding and Cake has $5 par, 60,000 shares outstanding

Choco Book Values Cake Book Values

Cake Fair Values

Cash and Receivable 350,000 180,000 170,000
Inventories 250,000 100,000 150,000
Land 700,000 120,000 240,000
Building and equipment 600,000 600,000 900,000
Patented technology 100,000 0 60,000
Accounts Payable 300,000 120,000 150,000
Long-term debt 0 400,000 350,000
Common Stock 750,000 300,000
Additional paid in capital 500,000 60,000
Retained earnings 12/31 450,000 120,000
Revenues 350,000 160,000
Expenses 310,000 130,000

Q1. How much is the consideration transferred?

Q2. What is the consolidated balance for Land?

Q3. What is the consolidated balance for Accounts payable?

Q4. Prepare fair value allocation and goodwill schedule at the date of the acquisition.
Answer #1

1.

Consideration = 40,000 X 32 = $ 1,280,000

2.

Consolidated balance of Land = 700,000 + 240,000 = $ 940,000

3.

Consolidated balance for Account Payable = 300,000 + 150,000 = $ 450,000

4.

Debit Credit
Cash and Receivable 170,000
Inventories 150,000
Land 240,000
Building and equipment 900,000
Patented technology 60,000
Goodwill 260,000
Accounts Payable 150,000
Long-term debt 350,000
Shareholder of Cake          1,280,000